OpenStudy (unklerhaukus):

\[\frac {\text d^2 x}{\text dt^2}=-\omega_0^2x+\beta x^2\]

OpenStudy (unklerhaukus):

\[\omega_0=2\pi\]

OpenStudy (anonymous):

i want to let \[ p=\frac{dx}{dt} \] then u have \[ \frac{d^2x}{dt^2}=\frac{dp}{dt}=\frac{dp}{dx}\frac{dx}{dt}=p \frac{dp}{dx} \] equation becomes to \( p \ dp=(-Ax+Bx^2) \ dx \) and \[ \ \frac{p^2}{2} =-\frac{A}{2}x^2+\frac{B}{3}x^2 +C\]

OpenStudy (anonymous):

and we arrive to \[ p^2=-Ax^2+\frac{2}{3}B x^3+C\]it gives \[(\frac{dx}{dt}) ^2=-Ax^2+\frac{2}{3}B x^3+C\] or \[\frac{dx}{\sqrt{-Ax^2+\frac{2}{3}B x^3+C}}=dt\]

OpenStudy (anonymous):

oh oh thats very bad for integration

OpenStudy (anonymous):

\[\frac {\text d^2 x}{\text dt^2}=-\omega_0^2x+\beta x^2\] \[ \int \text dt^2=\int \frac {\text d^2 x}{-\omega_0^2x+\beta x^2}\] \[ \text dt=[\ln (\frac { x}{-\omega_0^2+\beta x})+C \ ] \ dx\] \[\int \text dt=\int [\ln (\frac { x}{-\omega_0^2+\beta x})+C \ ] \ dx\] \[ \text t= x \ln (\frac { x}{-\omega_0^2+\beta x})+\frac{\omega_0^2}{\beta}\ln (-\omega_0^2+\beta x)+Cx+D \]

OpenStudy (anonymous):

i made a bunch of mistakes in integrating

OpenStudy (anonymous):

@eliassaab are we allowed to doing something like my second work?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

so thats wrong....and no need to correct my integration so what we gonna do !!!!

OpenStudy (unklerhaukus):

perhaps an analytic solution is impossible,

OpenStudy (anonymous):

yes; maybe

OpenStudy (experimentx):

seems that works http://www.wolframalpha.com/input/?i=y%27%27+%3D+ax+%2B+bx^2

OpenStudy (anonymous):

i think u made a little mistake santosh

OpenStudy (unklerhaukus):

ah @experimentX x=y t=x http://www.wolframalpha.com/input/?i=x%27%27+%3D+ax+%2B+bx%5E2

OpenStudy (experimentx):

ah ... my bad!!

OpenStudy (anonymous):

what a neat solution...lol

OpenStudy (experimentx):

that wasn't linear!!

OpenStudy (experimentx):

i guess the first method.

OpenStudy (anonymous):

it seems like motion equation but its extremely complicated

OpenStudy (experimentx):

acceleration depending on position?

OpenStudy (unklerhaukus):